OpenStudy (anonymous):

A square has a diagonal that is 50cm long. What is the area of the square? If necessary, round your answer to the nearest whole number.

OpenStudy (flying_dutchman):

 so, a = b as it's a square. 2500 = 2a^2 1250 = a^2 a = 35.35533906 area = a*a area = 1250